OpenStudy (anonymous):

21x(x+3)+7x^2(x+3) Factoring out the GCF.

OpenStudy (gorv):

(x+3)*(21x+7x^2)=(x+3)*7x(x+3)=7x*(x+3)^2

OpenStudy (anonymous):

why is it to the power of two ?

OpenStudy (gorv):

because x+3 is 2 times there

OpenStudy (gorv):

@Jlynn89

OpenStudy (anonymous):

I just had it written 7x(x+3)_ haha is why i was asking. I had it written like my other one where I did not do that.

OpenStudy (gorv):

no no they are in multiplication

OpenStudy (gorv):

@Jlynn89

OpenStudy (anonymous):

Thanks ^^; I understand why it's to the power of 2 now.

OpenStudy (gorv):

wlcme :)